OpenStudy (anonymous):

On Bill's football team, 1/3 of the players walk to practice and 25% are driven by their parents. The remaining 15 players take the bus. How many members are on the football team?

OpenStudy (rainbow_dash):

If 34 is an option, I say click that :3

OpenStudy (rainbow_dash):

I just did 15 * .75 * .33, which came up with 34....but hero knows what he's doing. i don't really..

hero (hero):

Hint: \(\large B - \frac{1}{3}B - \frac{1}{4}B = 15\)

hero (hero):

The more simplified version of that is: \(\large B - \frac{B}{3} - \frac{B}{4} = 15\)

hero (hero):

Now all you have to do is solve for B

OpenStudy (anonymous):

I'm really confused on how to solve for B

hero (hero):

Well, there are a series of steps you could to solve for B

hero (hero):

Now you can just combine fractions and cross multiply

OpenStudy (anonymous):

Combine which fractions?

hero (hero):

Correction: \(\large \frac{3B}{3} - \frac{B}{3} = \frac{60}{4} + \frac{B}{4}\)

hero (hero):

If you combine the fractions, you end up with \(\large \frac{2B}{3} = \frac{60 + B}{4} \)

hero (hero):

Now, just cross multiply and continue solving.

hero (hero):

If you can tell me what B equals, I'll give you a medal @jazzyrocroyal

OpenStudy (anonymous):

@Hero 36